# Multi-wavelength observations of the Luminous Fast Blue Optical Transient AT 2023fhn

### v1

@@ Abstract
Results. We find that the neither the host galaxy nor circumstellar environment of AT 2023fhn are unusual compared with previous LFBOTs, but that AT 2023fhn has a much lower X-ray to ultraviolet luminosity ratio than previous events.
@@
Since AT 2018cow, several more LFBOTs have been discovered. Confirmed events include AT 2018lug/ZTF 18abvkwla (‘the Koala’, Ho et al., 2020), CSS161010 (Coppejans et al., 2020), AT 2020xnd/ZTF 20 acigmel (‘the Camel’, Perley et al., 2021; Bright et al., 2022; Ho et al., 2022), AT 2020mrf (Yao et al., 2022), AT 2022tsd (‘the Tasmanian Devil’, Matthews et al., 2023) and AT 2023fhn (‘the Finch’, Chrimes et al., 2024). Despite variety (e.g. in peak luminosity), they share the same key features of hot, largely featureless spectra at early times, optical luminosities rivalling gamma-ray burst (GRB) afterglows and superluminous supernovae, plus bright X-ray and radio emission. They are estimated to occur at 0.1% of the local core-collapse supernova rate (Ho et al., 2023d).
@@
Several models have been put forward to explain LFBOTs. Tidal disruptions of compact, hydrogen-poor stars (such as white dwarfs) around intermediate mass black holes (IMBHs) can plausibly explain the optical rise and fall timescale, spectral features and X-ray variability timescale (Perley et al., 2019; Kuin et al., 2019). However, the dense CSM inferred from radio observations is hard to explain in such a scenario (e.g. Margutti et al., 2019). Other possibilities include failed supernovae, in which a black hole is formed and the emission is powered by accretion onto the natal black hole rather than radioactive decay in the ejecta (Perley et al., 2019; Quataert et al., 2019), choked jets (e.g. Gottlieb et al., 2022; Soker, 2022), highly aspherical supernovae (‘ellipsars’, DuPont et al., 2022), and the mergers of compact objects and/or massive stars (Lyutikov & Toonen, 2019; Uno & Maeda, 2020; Schrøder et al., 2020), such as black holes and Wolf-Rayet stars (Metzger, 2022). A dense outflow from the progenitor may result in dust echoes (Metzger & Perley, 2023). CSM shock interaction models have also been put forward (e.g. Fox & Smith, 2019; Xiang et al., 2021; Pellegrino et al., 2022; Khatami & Kasen, 2023), but the X-ray variability, broadband spectral evolution, late-time UV/X-ray emission from AT 2018cow and giant optical flares from AT 2022tsd all indicate the presence of a central engine.
@@ 3.1 Local environment
The second epoch of HST imaging presented in this paper allows us to examine the environment directly underlying the transient after it has faded. As noted by Chrimes et al. (2024), there is diffuse emission in the vicinity of the transient. To characterise this faint underlying population, we place 0.2 arcsec (and 0.4 arcsec) apertures at the location of AT 2023fhn in all six epoch 2 images. The images are aligned with - shifts using 5 common point sources in every image, with respect to the location of AT 2023fhn in the epoch 1 F555W image. The rms of these relative astrometric alignments is 5–10 mas, better than the absolute astrometry of the images (which have been aligned with the Gaia DR3 reference frame), and much smaller than the aperture size. We perform photometry with photutils, estimating the background with either the median image background (with medianbackground) or an annulus (1.5 to 4 times the aperture radius, with pixels values clipped at 3). The appropriate encircled energy corrections for each filter and aperture are applied. Magnitudes are then calculated using the photplam and photflam header keywords111https://hst-docs.stsci.edu/wfc3dhb/chapter-9-wfc3-data-analysis/9-1-photometry, and are listed in Table 4. The only detections are in F555W and F814W. To investigate the nature of these detections, we place eight 0.4 arcsec apertures at equal spacing around the location of AT 2023fhn in a circle of radius 20 pixels (0.5 arcsec). With the F555W filter and median background subtraction, we have significant detections in 5/8 apertures, with a mean magnitude of 25.90.6 in these apertures - consistent with the measurement at the precise location of AT 2023fhn. This demonstrates that the emission in this area is from an extended, diffuse background, rather than any significant contribution by residual light from AT 2023fhn. This can also be seen in Table 4, where the magnitudes calculated with annulus background subtraction are fainter, since the local background is elevated. Larger apertures also give brighter magnitudes, despite encircled energy correction (unlike point sources in the field). We similarly disfavour any significant contribution from a compact cluster at this specific location, which would appear as a point source in the image given the physical scale at this redshift of 100 pc pixel-1. However, the presence of a globular cluster (which would favour an IMBH interpretation, e.g., Lützgendorf et al., 2013) cannot be ruled out, as even the brightest globular clusters would be far below detection limits at this distance and limiting magnitude (Chrimes et al., 2024). Shifting the circle of apertures 5 arcsec to the north, well away from the galaxies, we find non-detections in all eight apertures with a 3 upper limit of 26.7. We therefore conclude that there is extended, diffuse emission from an underlying stellar population at the location of AT 2023fhn.
@@
We now estimate the age and dust extinction of this underlying population. First, we correct for the (low) Galactic extinction of =0.0254 (Schlafly & Finkbeiner, 2011)222https://irsa.ipac.caltech.edu/applications/DUST/ using the filter effective wavelengths (Rodrigo et al., 2012; Rodrigo & Solano, 2020) and the Python extinction package (Barbary, 2016) with a Fitzpatrick (1999) extinction law and . To estimate the age and local (intrinsic) extinction, we fit the Galactic-extinction corrected and photometry to BPASS (Binary Population and Spectral Synthesis v2.1, Eldridge et al., 2017; Stanway & Eldridge, 2018) single-age spectral templates. These are constructed by assuming that a stellar population of 106M⊙ is formed instantaneously, and left to evolve with no further star formation. We use these simple stellar populations since the limited data available to model solely the local environment of AT 2023fhn precludes a more complex procedure, including, for example, the star-formation history (however, see the next section). A fixed metallicity of half-Solar is adopted ( by mass fraction). We therefore simply fit for the age of the population, the luminosity (i.e. mass) of the stellar population is then allowed to freely vary to minimise . Four data points are used ( and ) where the upper limits are treated as data points with zero flux and an uncertainty equal to the flux of the 1 upper limit. We therefore have 2 fit parameters and 4 data points for 2 degrees of freedom. Fitting is performed by multiplying the (de-redshifted) filter response curves (Rodrigo et al., 2012; Rodrigo & Solano, 2020) with the BPASS spectra to extract fluxes and hence magnitudes from the spectra. These are compared with the absolute magnitudes in each filter, after correction for a range of intrinsic extinction values from =0.0 to 1.0. The intrinsic extinction correction uses the rest-frame effective wavelength of each filter. The and filters are not used in this fit since the upper limits are shallower than the and detections, and so provide no additional constraints.
The results are shown in Figure 2. The top panel shows the best-fit single-age BPASS spectrum. The lower panel shows log across the parameter space (indicated by the shading). Each pixel represents a unique combination of and a BPASS simple stellar population at a given age. The 68% and 90% confidence intervals are indicated by white contours (where the intervals are from Avni, 1976).
@@
We also measure the local surface brightness in epoch 2 (in a 0.5 arcsec radius around AT 2023fhn’s position), giving 25.1 mag arcsec-2 in and 24.65 mag arcsec-2 in . This compares well with the 25.2 mag arcsec-2 and 24.6 mag arcsec-2 values from the transient-subtracted images in Epoch 1 (see Chrimes et al., 2024). The surface brightness is 25.76 mag arcsec-2, which after Galactic extinction correction is 25.27 mag arcsec-2. The rest-frame central wavelength of is 2700Å. This allows for a better comparison with the UV () surface brightness distribution for supernova environments, as reported by Kelly & Kirshner (2012) than made by Chrimes et al. (2024) with . The Galactic extinction-corrected surface brightness is in the faintest 10% for local supernova values; this is therefore faint but not unprecedented. We note that supernovae type IIb are the most likely supernova sub-class to explode in young, but low surface brightness environments, and are also found at the highest host-normalised offsets on average (Kelly & Kirshner, 2012).
@@
![Refer to caption](extracted/5678772/region999_bestfit_SED.png)
![Refer to caption](extracted/5678772/chisq.png)
@@ 3.2 Global host properties
We attempt to collect as close to 100% of the galaxy light from HST photometry as possible. We measure the Petrosian radius (Petrosian, 1976) of the spiral galaxy with the statmorph package (Rodriguez-Gomez et al., 2019, ) and adopt 1.5 as a radius that encloses 100% of the flux (e.g. Conselice, 2003). We account for the projected ellipticity and orientation of the galaxy using the ellip and theta outputs. A pixel mask is produced using these parameters as measured from the image, and applied to the other HST images, as shown in Figure 3. The flux within the mask is summed, and background subtraction (as for the local environment measurements above) uses the sigma-clipped median background, scaled for the number of pixels in the mask. Repeating the procedure for the satellite galaxy produces a 1.5 pixel mask that lies entirely within the spiral’s mask. We therefore use spiral pixel mask alone as it captures 100% of the flux from both galaxies.
@@
To supplement the HST data we add host photometry from archival catalogues. For additional optical points we use PanSTARRS data release 2 (Chambers et al., 2016). We use the catalogued Kron magnitudes (Kron, 1980, in , , , and ), which capture 90% of the light of extended sources, and increase the fluxes by a further 10% to approximate the 100% flux value 333https://outerspace.stsci.edu/display/PANSTARRS/PS1+Kron+photometry+of+extended+sources. The Kron radii for the spiral (4.41, 4.62, 4.36, 3.33 and 2.89 arcsec in respectively) extend past the position of the satellite in , so the system can be considered blended. Effective wavelengths for these filters are from Tonry et al. (2012). We also add far-UV and near-UV photometry from GALEX (Martin et al., 2003), plus W1, W2 and W3 detections from WISE. The spiral and satellite cannot be separated at the spatial resolution of these surveys, and neither galaxy is detected in 2MASS. The full list of photometry used to performed SED fitting is provided in Table 5.
@@
To perform SED fitting we use prospector (Leja et al., 2017; Johnson et al., 2021), which makes use of FSPS (Flexible Stellar Population Synthesis Conroy et al., 2009; Conroy & Gunn, 2010) and Python-FSPS (Johnson et al., 2023). For the Markov Chain Monte Carlo (MCMC) implementation we use emcee (Foreman-Mackey et al., 2013). We again use BPASS (Binary Population and Spectral Synthesis v2.1, Eldridge et al., 2017; Stanway & Eldridge, 2018) for the spectral models. Before being passed to prospector, the input photometry is corrected for Galactic extinction (as described in Section 3.1). We fit four parameters: the stellar mass , intrinsic extinction , population age and the timescale for an exponentially declining star-formation history . The redshift is fixed at , and the luminosity distance at Mpc.
@@
We run the MCMC with 128 walkers and 512 iterations; the full list of MCMC set-up parameters and joint posterior distributions (in the form of a corner plot) are provided in Appendix A. The maximum a posterior (MAP) spectrum is shown in Figure 4, with the associated properties from the posterior distribution listed in Table 6. Thus far, the metallicity has been fixed at half-Solar, based on the approximate mass of M⊙ and the mass-metallicity relation (Tremonti et al., 2004; Gallazzi et al., 2005). A similar table containing the results when metallicity is allowed to vary is also provided in Appendix A. In this case, the mass and SFR are similar, such that fixing at a more realistic value does not change our results in a qualitative sense. In the delayed- model, the current star-formation rate (SFR) is proportional to . The absolute value is obtained by normalisation with respect to the mass formed, yielding a SFR of 4 M⊙ yr-1. The galaxy pair is therefore dominated by a fairly typical star-forming spiral, but is perhaps notable for the likely presence of tidal interactions between the spiral and its satellite. In Figure 5 we plot its mass versus SFR, comparing with the host galaxies of previous LFBOTs. The galaxy has a high SFR and mass for LFBOT hosts, lying slightly above average in terms of specific star formation rate (sSFR), but well below the sSFR of the host of ZTF 18abvkwla.
![Refer to caption](x2.png)

### v2

@@ Abstract
Results. We find that neither the host galaxy nor circumstellar environment of AT 2023fhn are unusual compared with previous LFBOTs, but that AT 2023fhn has a much lower X-ray to ultraviolet luminosity ratio than previous events.
@@
Since AT 2018cow, several more LFBOTs have been discovered. Confirmed events include AT 2018lug/ZTF 18abvkwla (‘the Koala’, Ho et al., 2020), CSS161010 (Coppejans et al., 2020), AT 2020xnd/ZTF 20 acigmel (‘the Camel’, Perley et al., 2021; Bright et al., 2022; Ho et al., 2022), AT 2020mrf (Yao et al., 2022), AT 2022tsd (‘the Tasmanian Devil’, Matthews et al., 2023) and AT 2023fhn (‘the Finch’, Chrimes et al., 2024). Despite variety (e.g. in peak luminosity), they share the same key features of hot, largely featureless spectra at early times, optical luminosities rivalling superluminous supernovae, plus bright X-ray and radio emission. They are estimated to occur at 0.1% of the local core-collapse supernova rate (Ho et al., 2023d).
@@
Several models have been put forward to explain LFBOTs. Tidal disruptions of compact, hydrogen-poor stars (such as white dwarfs) around intermediate mass black holes (IMBHs) can plausibly explain the optical rise and fall timescale, spectral features and X-ray variability timescale (Perley et al., 2019; Kuin et al., 2019). However, the dense CSM inferred from radio observations is hard to explain in such a scenario (e.g. Margutti et al., 2019). Other possibilities include failed supernovae, in which a black hole is formed and the emission is powered by accretion onto the natal black hole rather than radioactive decay in the ejecta (Perley et al., 2019; Quataert et al., 2019), choked jets (e.g. Gottlieb et al., 2022; Soker, 2022), highly aspherical supernovae (‘ellipsars’, DuPont et al., 2022), and the mergers of compact objects and/or massive stars (Lyutikov & Toonen, 2019; Uno & Maeda, 2020; Schrøder et al., 2020), such as black holes and Wolf-Rayet stars (Metzger, 2022). A dense outflow from the progenitor may result in dust echoes (Metzger & Perley, 2023). CSM shock interaction models have also been put forward (e.g. Fox & Smith, 2019; Xiang et al., 2021; Pellegrino et al., 2022; Khatami & Kasen, 2024), but the X-ray variability, broadband spectral evolution, late-time UV/X-ray emission from AT 2018cow and giant optical flares from AT 2022tsd all indicate the presence of a central engine.
@@ 3.1 Local environment
The second epoch of HST imaging presented in this paper allows us to examine the environment directly underlying the transient after it has faded. As noted by Chrimes et al. (2024), there is diffuse emission in the vicinity of the transient. To characterise this faint underlying population, we place 0.2 arcsec (and 0.4 arcsec) apertures at the location of AT 2023fhn in all six epoch 2 images. The images are aligned with - shifts using 5 common point sources in every image, with respect to the location of AT 2023fhn in the epoch 1 F555W image. The rms of these relative astrometric alignments is 5–10 mas, better than the absolute astrometry of the images (which have been aligned with the Gaia DR3 reference frame), and much smaller than the aperture size. We perform photometry with photutils, estimating the background with either the median image background (with medianbackground) or an annulus (1.5 to 4 times the aperture radius, with pixels values clipped at 3). The appropriate encircled energy corrections for each filter and aperture are applied. Magnitudes are then calculated using the photplam and photflam header keywords111https://hst-docs.stsci.edu/wfc3dhb/chapter-9-wfc3-data-analysis/9-1-photometry, and are listed in Table 4. The only detections are in F555W and F814W. To investigate the nature of these detections, we place eight 0.4 arcsec apertures at equal spacing around the location of AT 2023fhn in a circle of radius 20 pixels (0.5 arcsec). With the F555W filter and median background subtraction, we have significant detections in 5/8 apertures, with a mean magnitude of 25.90.6 in these apertures - consistent with the measurement at the precise location of AT 2023fhn. This demonstrates that the emission in this area is from an extended, diffuse background, rather than any significant contribution by residual light from AT 2023fhn. This can also be seen in Table 4, where the magnitudes calculated with annulus background subtraction are fainter, since the local background is elevated. Larger apertures also give brighter magnitudes, despite encircled energy correction (unlike point sources in the field). For this reason, we exclusively use median background subtracted measurements - which are not biased by the nearest pixels in the image - in the following analysis.
We disfavour a significant contribution from a compact cluster at this specific location, which would appear as a point source in the image given the physical scale at this redshift of 100 pc pixel-1. However, the presence of a globular cluster (which would favour an IMBH interpretation, e.g., Lützgendorf et al., 2013) cannot be ruled out, as even the brightest globular clusters would be far below detection limits at this distance and limiting magnitude (Chrimes et al., 2024). Shifting the circle of apertures 5 arcsec to the north, well away from the galaxies, we find non-detections in all eight apertures with a 3 upper limit of 26.7. We therefore conclude that there is extended, diffuse emission from an underlying stellar population at the location of AT 2023fhn.
@@
We now estimate the age and dust extinction of this underlying population. First, we correct for the (low) Galactic extinction of =0.0254 (Schlafly & Finkbeiner, 2011)222https://irsa.ipac.caltech.edu/applications/DUST/ using the filter effective wavelengths (Rodrigo et al., 2012; Rodrigo & Solano, 2020) and the Python extinction package (Barbary, 2016) with a Fitzpatrick (1999) extinction law and . To estimate the age and local (intrinsic) extinction, we fit the Galactic-extinction corrected and photometry (both 0.2 and 0.4 arcsec apertures, median background subtracted) to BPASS (Binary Population and Spectral Synthesis v2.1, Eldridge et al., 2017; Stanway & Eldridge, 2018) single-age spectral templates. These are constructed by assuming that a stellar population of 106M⊙ is formed instantaneously, and left to evolve with no further star formation. We use these simple stellar populations since the limited data available to model solely the local environment of AT 2023fhn precludes a more complex procedure, including, for example, the star-formation history (however, see the next section). A fixed metallicity of half-Solar is adopted ( by mass fraction). We therefore simply fit for the age of the population, the luminosity (i.e. mass) of the stellar population is then allowed to freely vary to minimise . Four data points are used ( and ). Formal flux measurements are used for and ; these produce magnitudes of 29.0 and 27.3 respectively (with a 0.2 arcsec aperture) and 26.9 and 26.7 (with a 0.4 arcsec aperture). We therefore have 2 fit parameters and 4 data points for 2 degrees of freedom. Fitting is performed by multiplying the (de-redshifted) filter response curves (Rodrigo et al., 2012; Rodrigo & Solano, 2020) with the BPASS spectra to extract fluxes and hence magnitudes from the spectra. These are compared with the absolute magnitudes in each filter, after correction for a range of intrinsic extinction values from =0.0 to 4.0. The intrinsic extinction correction uses the rest-frame effective wavelength of each filter. The and filters are not used in this fit since the upper limits are shallower than the and detections, and so provide no additional constraints.
The results are shown in Figure 2. The left-hand panels show the best-fit single-age BPASS spectra. The right-hand panels show log across the parameter space. Each pixel represents a unique combination of and a BPASS simple stellar population at a given age. The 68% and 90% confidence intervals are indicated by crosses and dots respectively (where the intervals are from Avni, 1976). We can see that the choice of a 0.2 or 0.4 arcsec aperture makes little difference, likely because the increased aperture size simply captures more of the same diffuse background light at that position, without changing the colours significantly. In both cases, there is a degeneracy between age and extinction, and while the uncertainties are large, population ages in excess of a few hundred Myr are disfavoured. The range of possible ages for this diffuse emission is therefore broadly consistent with the core-collapse supernova delay time distribution (e.g. Eldridge et al., 2019).
@@
We also measure the local surface brightness in epoch 2 (in a 0.5 arcsec radius around AT 2023fhn’s position), giving 25.1 mag arcsec-2 in and 24.65 mag arcsec-2 in . This compares well with the 25.2 mag arcsec-2 and 24.6 mag arcsec-2 values from the transient-subtracted images in Epoch 1 (see Chrimes et al., 2024). The surface brightness is 25.76 mag arcsec-2, which after Galactic extinction correction is 25.27 mag arcsec-2. The rest-frame central wavelength of is 2700Å. This allows for a better comparison with the UV () surface brightness distribution for supernova environments, as reported by Kelly & Kirshner (2012) than made by Chrimes et al. (2024) with . The Galactic extinction-corrected surface brightness is in the faintest 10% for local supernova values; this is therefore faint but not unprecedented.
@@ 3.2 Global host properties
We attempt to collect as close to 100% of the galaxy light from HST photometry as possible. We measure the Petrosian radius (Petrosian, 1976) of the spiral galaxy with the statmorph package (Rodriguez-Gomez et al., 2019, ) and adopt 1.5 as a radius that encloses 100% of the flux (e.g. Conselice, 2003). We account for the projected ellipticity and orientation of the galaxy using the ellip and theta outputs. A pixel mask is produced using these parameters as measured from the image, and applied to the other HST images, as shown in Figure 3. The flux within the mask is summed, and background subtraction (as for the local environment measurements above) uses the sigma-clipped median background, scaled for the number of pixels in the mask. Repeating the procedure for the satellite galaxy produces a 1.5 pixel mask that lies entirely within the spiral’s mask. We therefore use spiral pixel mask alone as it captures 100% of the flux from both galaxies. Uncertainties are determined using the standard deviation of background pixels outside the galaxy mask, again using background estimator, the total background noise on the measurement is then . To investigate the effect of background variations, we split each image into regions of pixels - the approximate size of the galaxy mask - and determine the median background level in each region. The standard deviation of these, i.e. the uncertainty in the background level, is added in quadrature as an additional source of error on our HST galaxy photometry.
@@
To supplement the HST data we add host photometry from archival catalogues. For additional optical points we use PanSTARRS data release 2 (Chambers et al., 2016). We use the catalogued Kron magnitudes (Kron, 1980, in , , , and ), which capture 90% of the light of extended sources, and increase the fluxes by a further 10% to approximate the 100% flux value 333https://outerspace.stsci.edu/display/PANSTARRS/PS1+Kron+photometry+of+extended+sources. The Kron radii for the spiral (4.41, 4.62, 4.36, 3.33 and 2.89 arcsec in respectively) extend past the position of the satellite in , so the system can be considered blended in these filters. In and this is not the case, so they likely underestimate the flux from the combined galaxy-satellite system, and indeed these points lie a few sigma below the best-fit spectrum when they are included in the fit. We therefore remove the and filters from the fitting. We also add far-UV and near-UV photometry from GALEX (Martin et al., 2003), plus W1, W2 and W3 detections from WISE. Effective wavelengths for these filters are taken from Tonry et al. (2012). The spiral and satellite cannot be separated at the spatial resolution of these surveys, and neither galaxy is detected in 2MASS. The full list of photometry used to performed SED fitting is provided in Table 5.
@@
To perform SED fitting we use prospector (Leja et al., 2017; Johnson et al., 2021), which makes use of FSPS (Flexible Stellar Population Synthesis Conroy et al., 2009; Conroy & Gunn, 2010) and Python-FSPS (Johnson et al., 2023). For the Markov Chain Monte Carlo (MCMC) implementation we use emcee (Foreman-Mackey et al., 2013). We again use BPASS (Binary Population and Spectral Synthesis v2.1, Eldridge et al., 2017; Stanway & Eldridge, 2018) for the spectral models. Before being passed to prospector, the input photometry is corrected for Galactic extinction (as described in Section 3.1). We fit four parameters: the stellar mass , intrinsic (local to the transient) extinction , population age and the timescale for an exponentially declining star-formation history . The redshift is fixed at , and the luminosity distance at Mpc.
@@
We run the MCMC with 128 walkers and 512 iterations; the full list of MCMC set-up parameters and joint posterior distributions (in the form of a corner plot) are provided in Appendix A. The maximum a posterior (MAP) spectrum is shown in Figure 4, with the associated properties from the posterior distribution listed in Table 6. Thus far, the metallicity has been fixed at half-Solar, based on the approximate mass of M⊙ and the mass-metallicity relation (Tremonti et al., 2004; Gallazzi et al., 2005). A similar table containing the results when metallicity is allowed to vary is also provided in Appendix A. In this case, the mass and SFR are similar, such that fixing at a more realistic value does not change our results in a qualitative sense. In the delayed- model, the current star-formation rate (SFR) is proportional to . The absolute value is obtained by normalisation with respect to the mass formed, yielding a SFR of 5.5 M⊙ yr-1. In Figure 4, the F845M point lies above the best fit spectrum, possibly due to the presence of H emission in this medium-width filter. Using the offset between the F845M point and the best-fit spectrum, and assuming the offset is dominated by H emission, we estimate the line flux as 1erg s-1 cm-2 , which corresponds to a SFR of 1 M⊙yr-1 (Kennicutt et al., 1994). However, since the F845M point is only offset by 3, and given the uncertainty in the continuum level, we use continue to use the prospector derived SFR as the fiducial value.
In summary, the galaxy pair is dominated by a fairly typical star-forming spiral, but is perhaps notable for the likely presence of tidal interactions between the spiral and its satellite. In Figure 5 we plot its mass versus SFR, comparing with the host galaxies of previous LFBOTs. The galaxy has a high SFR and mass for LFBOT hosts, lying slightly above average in terms of specific star formation rate (sSFR), but well below the sSFR of the host of ZTF 18abvkwla.